a farmer was asked how many cows he had on his property. he replied that he was unsure, but he knew that when he counted them by twos, threes, fours, fives, or sixes, he always had one left over. the only way he could avoid this was to count by sevens; he then had none left over. what is the smallest number of cows the farmer could possibly have owned?
Answer by ankor@dixie-net.com(22740)

a farmer was asked how many cows he had on his property.
he replied that he was unsure, but he knew that when he counted them by twos, threes, fours, fives, or sixes, he always had one left over.
the only way he could avoid this was to count by sevens; he then had none left over.
what is the smallest number of cows the farmer could possibly have owned?
:
We know it is a multiple of 7, that it is an odd number, has to end in 1
multiply by numbers ending in 3 to get a 1
13 * 7 = 91, no
23 * 7 = 161, no
33 * 7 = 231, no
43 * 7 = 301, looks like a winner